At 24° latitude, the gravitational acceleration as a function of elevation z above sea level is given by g=a-bz,


where a=9.807 m/s2 and b=3.32x10-6 s-2. Determine the height above sea where the weight of an object will decrease by 1 percent.

1
Expert's answer
2021-11-08T08:49:41-0500

At a height above sea level, weight of object decreases by 1%

"\frac{g_o-g}{g_o}=\frac{1}{100}"

Given g=a-bz

At z=0, go=a-b(0)=a

go=9.807m/s2

"\frac{9.807-g}{9.807}=\frac{1}{100}=0.01"

9.807-g=0.01*9.807

9.807-g=0.09807

g=9.807-0.09807

=9.70893m/s2

g=a-bz

"z=\frac{a-g}{b}"

"=\frac{9.807-9.70893}{3.32*10^{-6}}"

=29539.156m